Indicate which of the following sets of lines are parallel, perpendicular, or neither. 3x + y = 5. X + y = 3.
Answer by algebrahouse.com(1659)

Parallel lines have the same slope.
Perpendicular lines have slopes which are negative reciprocals.

Slope-intercept form is y = mx + b
m is the slope of the line
b is the y-intercept of the line

3x + y = 5
y = -3x + 5 {subtracted 3x from both sides}
slope = -3

x + y = 3
y = -x + 3 {subtracted x from both sides}
slope = -1

= neither
These two lines are neither, parallel nor perpendicular, because
the slopes are not equal, nor are they negative reciprocals.

For more help from me, visit: www.algebrahouse.com